Graph the function fix) 2- for x 2 1 - 3x, for x 2 OA OB Ос. OD Click to select your answer 30 004
Answer #1

1-32 f(n) = 2-2 for xsz for x 2 The meaning of above data if we take upto a less than 37 eoqual to 2 then Consider the functiif X = 2 y flul=1-32=1-3(2)=1-6=-5 But value not included in line. hole. indicate who with x=4 9= 1-3(4) = 1-12=-|| 10 8 6 4